### v1

# Superradiant Bosons Driving Supermassive Black Hole Mergers

Abstract
Ultralight bosons (ULBs) can form macroscopic superradiant clouds around spinning black holes. We show that for scalar ULB masses $\mu\sim 10^{-22}-10^{-21}$ eV boson cloud dynamical friction drives supermassive black hole (SMBH) final-parsec evolution in $\lesssim 1$ Gyr and suppresses the nanohertz gravitational wave background with turnover. Considering century-monitored OJ287 system, we place novel bounds restricting ULB mass range $\mu\simeq(8.5-22)\times 10^{-22}$ eV independent of any dark matter assumptions and show ULB drag can also efficiently reconcile any future confirmations of the debated orbital decay excess. Forthcoming pulsar timing array data and precise SMBH orbital timings will decisively test this scenario.
preprint: CTPU-PTC-25-18, KEK-QUP-2025-0010, KEK-TH-2719, KEK-Cosmo-0380
Introduction.— Ultralight bosons (ULBs) – hypothetical quantum fields with masses below the eV scale – naturally arises in many extensions of the Standard Model (SM). Prominent examples include the QCD axion Peccei:1977hh ; Weinberg:1977ma ; Wilczek:1977pj ; Arvanitaki:2009fg , “fuzzy” dark matter (DM) candidates with masses $\mu\sim 10^{-22}$ eV Turner:1983he ; Press:1989id ; Sin:1992bg ; Hu:2000ke ; Goodman:2000tg , and majoron or axion-like fields Chikashige:1980ui ; Gelmini:1980re ; Arvanitaki:2009fg . Their small masses can be ensured by approximate global symmetries broken at high energies. With feeble ULB non-gravitational couplings, cosmological and astrophysical probes (see, e.g. Kimball:2023vxk ) can be especially sensitive beyond and complementary to laboratory searches.
Rapidly rotating (Kerr) black holes (BHs) can efficiently amplify bosonic field waves, extracting rotational energy and forming long-lived macroscopic “gravitational atom” (GA) condensates Arvanitaki:2009fg ; Arvanitaki:2010sy ; Brito:2015oca ; Baumann:2019eav . This superradiant amplification is gravitational analogue of observed Dicke superradiance in quantum optics and condensed matter Dicke:1954zz . The resulting boson clouds have rich phenomena, including monochromatic gravitational wave (GW) emission yoshino2014gravitational ; Chan:2022dkt , restricting population of spinning BHs Ng:2020ruv ; Stott:2020gjj ; Davoudiasl:2019nlo as well as inducing level-mixing and ionization transitions in binaries Baumann:2019ztm ; Ding:2020bnl ; Tomaselli:2023ysb . This offers an intriguing window for exploring fundamental physics.
Supermassive BHs (SMBHs) constitute prime timely targets for investigation. Recent James Webb Space Telescope (JWST) observations have uncovered active galactic nuclei at high redshifts of $z\simeq 10$ Matthee:2023utn ; 2023Natur.621…51D , implying rapid SMBH growth, while pulsar-timing arrays (PTAs) report detection of a stochastic GW background in the nanohertz band NANOGrav:2023gor ; EPTA:2023fyk ; Zic:2023gta ; Xu:2023wog . These observations sharpen the long-standing “final-parsec” puzzle, how SMBH binaries efficiently shrink on sub-pc scales in Hubble time Milosavljevic:2002ht .
In this work, we demonstrate that superradiant ULBs can significantly affect SMBH binary dynamics and efficiently drive merger evolution in the final parsec regime. An excellent testbed is offered by an archetypal SMBH binary OJ287 already monitored for over a century. Recent analysis has claimed potential excess in its orbital-decay power 2024ApJ…962L..40C , although its existence remains debated Deb:2025raq . We derive novel limits on ULBs from OJ287 independent of cosmological abundance, and also show superradiant ULBs can readily reconcile discrepancy in its orbital dynamics if confirmed. We discuss implications for GWs and identify distinct novel signatures that can definitively test this scenario. We adopt $c=\hbar=G=1$.
Ultralight boson superradiance.— ULB fields can efficiently seed macroscopic superradiant clouds around spinning BHs when their Compton wavelength $\lambda_{c}=1/\mu$, with $\mu$ being their mass, exceeds Schwarzschild radius $r_{s}=2M_{\rm BH}$ of BH with mass $M_{\rm BH}$. As we show, superradiance clouds can significantly impact SMBH binary merger dynamics.
Let us consider a minimal scenario of massive scalar field $\Phi$ propagating on a Kerr spacetime background. with Lagrangian contributions $\mathcal{L}\supset-\left(\partial\Phi\right)^{2}/2-\mu^{2}\Phi^{2}/2$, neglecting possible self-interactions (see Arvanitaki:2010sy ; Baryakhtar:2020gao ; Omiya:2022gwu ; Chia:2022udn ; Xie:2025npy for their potential effects). The resulting Klein-Gordon equation governing the ULB behavior can be simplified to a Schrödinger-like equation by expanding it with respect to a small “gravitational fine structure constant” $\alpha=M_{\rm BH}\mu<1$. See Supplemental Material for details. In this regime, the Schwarzschild radius of BH is smaller than the Compton wavelength of ULB and the system can be described as “gravitational atom (GA)” resembling eigenfunctions of a hydrogen atom described by quantum numbers $n$, $l$, $m$ as
$\psi_{nlm}(r,\theta,\phi)\simeq R_{nl}(r)Y_{lm}(\theta,\phi)e^{-i(\omega_{nlm}-\mu)t}~{},$ (1)
where $R_{nl}(r)$ is the radial wave function and $Y_{lm}(\theta,\phi)$ are the spherical harmonics. The corresponding eigenvalue consists of a real part and an imaginary part $\omega_{nlm}=E_{nlm}+i\Gamma_{nlm}$, which denotes the energy of the eigenstate and its growth or decay rate respectively. Exponential superradiance growth occurs when $\Gamma_{nlm}>0$.
The superradiant growth rate of an $|nlm\rangle$ state can be computed analytically using the Detweiler approximation Detweiler:1980uk , considering $\alpha\lesssim 0.3$ Brito2015b ; Cannizzaro:2023jle , $\Gamma_{nlm}=2\,\tilde{r}_{+}\,C_{nl}\,g_{lm}(m\,\Omega_{\rm BH}-\omega_{nlm})\,\alpha^{4l+5}$, where $\Omega_{\rm BH}=\chi/2M_{\rm BH}\tilde{r}_{+}$ is the angular velocity of the outer BH horizon at $\tilde{r}_{+}=1+\sqrt{1-\chi^{2}}$ and the dimensionless BH Kerr spin parameter is defined as $\chi=J/M_{\rm BH}^{2}$, where $J$ being the angular momentum. Here, $C_{nl}$ are numerical coefficients and $g_{lm}=g_{lm}(\chi,\alpha,\mu)$ are functions Detweiler:1980uk ; Baumann:2019eav . In a superradiant state with $\Gamma_{nlm}>0$, ULB cloud will exponentially grow until spin $\chi$ becomes insufficient and saturates at $\chi_{\text{sat}}=4m\alpha/(m^{2}+4\alpha^{2})$.
By analogy with electric case, GA Bohr radius is $r_{0}=1/\mu\alpha$, and the excited states have radii $r_{n}\simeq n^{2}r_{0}$. The density of a boson cloud $\rho_{nlm}=M_{nlm}|\psi_{nlm}|^{2}$ can be described as (e.g. Cao:2023fyv )
$\rho_{nlm}(x,\theta)=$ $~{}2.72\times 10^{12}\,\mathrm{M_{\odot}\,pc^{-3}}A_{nlm}(x,\theta)\frac{\beta}{0.1}\left(\frac{\alpha}{0.1}\right)^{6}$
×$\times\left(\frac{1.83\times 10^{10}M_{\odot}}{M_{\rm BH}}\right)^{2}~{},$ (2)
where $A_{nlm}(x,\theta)$ denotes the spatial density distribution, where $x=r/r_{0}$ and $\beta$ is the mass ratio between the boson cloud and BH.
OJ287 dynamics.— An intriguing SMBH system to explore the dynamical effects of superradiance clouds is blazar OJ287. We illustrate our scenario in Fig. LABEL:fig:illustration. OJ287 is an unusually massive and tight BH binary with primary BH mass of $M_{\rm BH}=1.83\times 10^{10}\,M_{\odot}$, secondary BH mass of $M_{\ast}=1.50\times 10^{8}\,M_{\odot}$ and orbital timing measurements collected over $\sim 120$ yr period. Its orbital parameters include semi-major axis $a=0.056\,\mathrm{pc}$, inclination angle $\iota=90^{\circ}$, and eccentricity $e=0.657$ Dey:2018mjg ; 1997ApJ…481L…5V ; Valtonen:2013spa (see also Supplemental Material). The binary is shrinking due to continuous GW emission, expected from general relativity Peters:1964zz
$\langle P_{\rm GW}\rangle=-\dfrac{32m_{r}^{2}M_{t}^{3}}{5a^{5}}(1-e^{2})^{-7/2}\Big{(}1+\frac{73}{24}e^{2}+\dfrac{37}{96}e^{4}\Big{)}~{},$ (3)
where $m_{r}=M_{\rm BH}M_{\ast}/(M_{\ast}+M_{\rm BH})$ is the reduced system mass and $M_{t}=M_{\ast}+M_{\rm BH}$ is the total mass.
Recently, it has been reported that OJ287 apparent total orbital-decay power $\langle P_{\rm tot}\rangle=(3.66\pm 0.24)\times 10^{41}\,$ seemingly exceeds the quadrupole GW prediction $\langle P_{\rm GW}\rangle=(2.62\pm 0.02)\times 10^{41}\,$ by $\sim 4.3\sigma$ 2024ApJ…962L..40C . Notably, these estimates rely on period $T$ and $\dot{T}$ obtained from a GW-driven model. While this has been challenged in Deb:2025raq , including use of redshifted period, the suggested alternative approach of estimating $\langle P_{\rm tot}\rangle$, $T$ and $\dot{T}$ itself relies on using a common GW formalism. We set novel ULB limits from OJ287 dynamics and show that superradiant ULBs can account for the potential excesses in orbital decay if confirmed.
ULBs of mass $\mathcal{O}(10^{-21})\,\mathrm{eV}$ can form a superradiance cloud around primary SMBH of OJ287 within the current age of the Universe $\tau_{\mathrm{U}}\simeq 13.8\,\mathrm{Gyr}$, considering high initial spin that we assume for simplicity to be $\chi_{\rm ini}=1$ since the spins of most SMBHs are expected to be significant Volonteri:2004cf . This can be seen from the superradiance growth rate of the fastest growing GA state $|211\rangle$
$\Gamma_{211}\simeq 1.2\times 10^{-16}\left(\frac{\alpha}{0.1}\right)^{9}\left(\frac{1.8\times 10^{10}M_{\odot}}{M_{\rm BH}}\right)\,{\rm s}^{-1}~{}.$ (4)
The $|211\rangle$ cloud has a typical Bohr radius of $r_{0}\simeq\mathcal{O}(0.01-1)\,\mathrm{pc}$ and density of $\rho_{211}=M_{211}|\psi_{211}|^{2}$. Hence, $|211\rangle$ GA state can be fully saturated and with the cloud mass of $M_{211}\simeq\alpha M_{\rm BH}$ Tsukada:2018mbp . We note that the total cloud growth time takes number of e-folds that can be up to $\log(\Delta\chi M_{\text{BH}}^{2})\sim\mathcal{O}(150)$ times larger than $1/\Gamma_{211}$, which is still significantly less than $\tau_{U}$ for OJ287.
For superradiant GA cloud to prominently influence dynamics of SMBH binary, its configuration must be sufficiently stable against depletion on relevant timescales. For OJ287 we can estimate that the monochromatic GW emission Brito:2014wla ; Yoshino:2013ofa rate for the dominant $|211\rangle$ level is $\Gamma_{\mathrm{GW}}\simeq-2.11\times 10^{-22}\,\mathrm{s}^{-1}$. Hence, the respective depletion time exceeds the age of the Universe $\tau_{\mathrm{U}}$. In a binary system, Landau-Zener bound-to-bound resonances can occur when the companion’s orbital frequency matches the cloud level spacing Baumann:2019ztm ; Ding:2020bnl ; Takahashi:2021yhy ; Tong:2021whq ; Takahashi:2023flk . Considering the measured mass ratio $q=M_{\ast}/M_{\text{BH}}\simeq 8\times 10^{-3}$ and coupling $\alpha\simeq 0.12$, the survival fraction of the $|211\rangle$ state after the strongest $m=1$ crossing exceeds $99\%$ Baumann:2019ztm . Bound-to-free “ionization” resonances Baumann:2021fkf ; Tomaselli:2023ysb can eject bosons and deplete the cloud, which, however, is subdominant compared with the accretion effect of the companion for a massive cloud such as the case we consider here Baumann:2021fkf . A rough estimate of the timescale for accretion to consume the cloud is $\gtrsim 10^{8}{\rm yr}$, which is much larger than the duration of accretion experienced by the companion so far. We also estimate that tidal effects caused by the companion Cardoso:2020hca result in a small tidal parameter $\epsilon\sim 10^{-8}$ for OJ287, indicating stability to them. Hence, dominant depletion channels are insignificant, and GA cloud can be considered as quasi-static, consistent with our analysis. See Supplemental Material for additional details.
As the secondary BH of mass $M_{\ast}$ traverses superradiance GA cloud it experiences dynamical friction with dissipation power Hui:2016ltb
$P_{\mathrm{DF}}=\frac{4\pi q^{2}M_{\rm BH}^{2}}{v}\rho_{nlm}(x_{\ast},\theta_{\ast})C_{\Lambda}~{},$ (5)
where $x_{\ast}\equiv R_{\ast}/r_{0}$ and $\theta_{\ast}$ represents the angle between the BH spin direction and the companion location, $v$ is speed relative to the boson cloud that equals $\alpha\sqrt{(1+q)/x_{\ast}}$ in circular orbit and $\alpha\sqrt{(1+q)(2/x_{\ast}-r_{0}/a)}$ in elliptic orbit. Here, $C_{\Lambda}$ is the Coulomb logarithm that can be estimated as $C_{\Lambda}(y)=\mathrm{Cin}(2y)+\sin 2y/2y-1$. Since boson cloud density that the secondary BH pass during its orbit is varying, an effective dynamical friction power can be estimated by averaging contributions over the entire orbit as $P_{\mathrm{DF}}\right>=(1/T)\int_{0}^{2\pi}P_{\mathrm{DF}}(\varphi_{\ast})(\mathrm{d}t/\mathrm{d}\varphi_{\ast})\mathrm{d}\varphi_{\ast}$ where $\varphi_{\ast}$ and $T$ are the angular phase of the companion orbit with respect to the primary BH and the orbital period, respectively. GA back-reaction was shown to be well described as dynamical friction Tomaselli:2023ysb .
Figure 2: Orbital-decay power from a saturated GA superradiance cloud $|211\rangle$ state in OJ287, considering cloud mass fraction $\beta\simeq\alpha$. Novel upper limits considering ratio of $\langle P_{\rm DF}\rangle=\xi\langle P_{\rm GW}\rangle$ with $\xi=0.10,0.05,0.01$ (dashed, dot-dashed, dotted lines) between dynamical friction and GW radiation are displayed. Dissipation power $\langle P_{\rm tot}\rangle-\langle P_{\rm GW}\rangle$ (gray band) that can account for the claimed excess 2024ApJ…962L..40C . GA couplings where saturated spin $\chi_{211,{\rm sat}}$ is below the two measured primary values $\chi_{P_{1}}=0.313$ Valtonen:2016awd (dark blue region) and $\chi_{P_{2}}=0.381$ Dey:2018mjg (light blue region), parameters where $170\times\Gamma_{211}^{-1}>\tau_{\rm U}$ and cloud won’t form (red region) are shown.
[…]
We note that our scenario is also consistent with timing observations of OJ287 flares, which are thought to originate from secondary BH interacting with the primary BH accretion disk. Integrating the density profile in Eq. (Superradiant Bosons Driving Supermassive Black Hole Mergers) for the $|211\rangle$ state out to $50r_{s}$ yields a cloud mass of $\sim 0.05\%$ of the primary BH mass (see also Supplemental Material). This is well below the $\sim 3\%$ mass ratio at which the flare timing predictions could be impacted, as suggested by analysis of Alachkar:2022qdt .
We set novel limits on ULBs independent of assumptions on their cosmological abundance considering no excess of orbital dynamics over relativity predictions at $0.41\sigma$ at 1PN level Deb:2025raq . In Fig. 2 we show effects of different power ratio of $\langle P_{\rm DF}\rangle=\xi\langle P_{\rm GW}\rangle$ with $\xi=0.10,0.05,0.01$ between dynamical friction and GW radiation that enables constraining the boson mass yielding novel upper limits of $\mu<7.9,7.4,6.4\times 10^{-22}$ eV (dashed, dot-dashed, dotted curves). Observations limit power ratio $\langle P_{\rm DF}\rangle/\langle P_{\rm GW}\rangle\lesssim 0.01$ as demonstrated by Deb:2025raq . Hence, this forbids ULB masses larger than $6.4\times 10^{-22}\,\text{eV}$ considering superradiant effects. Combining with requirements that superradiant boson cloud formation satisfies $170\times\Gamma_{211}^{-1}<\tau_{\mathrm{U}}$, $\chi<\chi_{211,{\rm sat}}$, and Detweiler approximation $\alpha\lesssim 0.3$, we set novel bounds restricting ULB mass range $\mu\simeq(8.5-22)\times 10^{-22}$ eV.
Figure 3: Allowed SMBH mass and ULB mass parameter space where binary evolution time from $2\,\mathrm{pc}$ to $0.01\,\mathrm{pc}$ is shorter than $\tau_{\rm pc}<1\,\mathrm{Gyr}$, alleviating the final-parsec problem. The gray dashed line corresponds to conservative superradiance timescales of $170\times\Gamma_{211}^{-1}<\tau_{U}$. The parameters for addressing OJ287 orbital decay $(1.83\times 10^{10}\,M_{\odot},0.9\times 10^{-21}\,\mathrm{eV})$ are shown by a blue star.
Final parsec evolution.— GA superradiance clouds stemming from ULBs of mass around $\mu\simeq\mathcal{O}(10^{-21})\,\text{eV}$, as suggested by OJ287 orbital decay analysis, envelop $\mathcal{O}(10^{10})\,M_{\odot}$ SMBHs with Bohr radii $r_{c}\sim\mathcal{O}(1)$ pc. At this regime the evolution of SMBH binaries could stall, known as the “final-parsec problem” Milosavljevic:2002ht . Hence, the dynamical friction from GA boson clouds in SMBH binaries could efficiently accelerate the shrinkage of binary separation within final pc evolution.
We estimate that bound-to-bound resonance transitions deplete at most $\mathcal{O}(10)\%$ of the $|211\rangle$ population for equal-mass binaries and $\lesssim 10\%$ for $q=0.1$. See Supplemental Material for further details. This is distinct from OJ287 where GA is taken to be almost purely in $|211\rangle$ state. To capture this behavior we model the GA cloud distribution as $\rho=\eta\rho_{211}+(1-\eta)\rho_{31{-}1}$, with $\eta$ denoting relative contribution of $|211\rangle$ and $|31-1\rangle$ states that depends on SMBH binary parameters. For $q=0.1$ binaries we have $\eta\simeq 0.9$, while for $q=1$ case $\eta\simeq 0.1$. Here, we estimate negligible contributions of other states, although our approach can be readily extended to incorporate this. Then the orbital decay rate of SMBH binaries can be found by semi-major axis evolution as
[…]
The resulting elapsed time within final parsec $\tau_{\rm pc}$ can be estimated by integrating Eq. (6) as $\tau_{\rm pc}=\int qM_{\rm BH}^{2}/(2a^{2}\langle P_{\mathrm{tot}}\rangle)\mathrm{d}a$ over final parsec separation.
[…]
Figure 4: Characteristic GW background strain resulting from SMBH binaries considering ULB cloud dynamical friction effects for a range of boson masses. Shaded region above the blue solid line depicts all SMBHs having maximal initial spin, and shaded region above blue dashed line depicts conservative case of $170\times\Gamma_{211}^{-1}<\tau_{U}$. The vertical lines show turnover frequencies $f_{b}$ for different $(\mu,M_{\rm BH})$. Observations from NANOGrav NANOGrav:2023gor , EPTA EPTA:2023xxk , and PPTA Reardon:2023gzh are shown.
Gravitational wave background.— Once a SMBH binary evolution has crossed the final parsec, GW emission dominates and sources nanohertz stochastic GW background observable by PTAs. A superradiant cloud drag reduces the energy radiated per logarithmic frequency interval,
$\frac{\mathrm{d}E_{\mathrm{GW}}}{\mathrm{d}f_{s}}=\frac{\pi^{2/3}}{3}\frac{M_{\rm BH}^{5/3}}{f_{s}^{1/3}}\frac{q}{(1+q)^{1/3}}\frac{\langle P_{\mathrm{GW}}\rangle}{\langle P_{\mathrm{GW}}\rangle+\langle P_{\mathrm{DF}}\rangle}~{},$ (7)
with a characteristic strain
$h_{c}^{2}(f)=\frac{4}{\pi f}\int\mathrm{d}z\mathrm{d}q\mathrm{d}M_{\rm BH}\frac{\mathrm{d}^{3}n}{\mathrm{d}z\mathrm{d}q\mathrm{d}M_{\rm BH}}\frac{\mathrm{d}E_{\mathrm{GW}}}{\mathrm{d}f_{s}}~{}.$ (8)
where $f_{s}=f(1+z)$ is the source frequency and $\mathrm{d}^{3}n/\mathrm{d}z\mathrm{d}q\mathrm{d}M_{\rm BH}$ is the SMBH binary population that we estimate considering galaxy merger rate distribution, where we follow NANOGrav:2023hfp ; Alonso-Alvarez:2024gdz to evaluate SMBH population (see Supplemental Material). In evaluating Eq. (7) we account for distinct GA configurations by appropriately considering cloud distribution profile $\rho$ with ratio of states $\eta$ set depending on SMBH binary and ULB parameters. With superradiant growth relevant when $\Gamma_{211}^{-1}<\tau_{\mathrm{U}}$, we approximate that only SMBHs above a critical mass $M_{\rm crit}(\mu)$ host GA clouds. For our benchmark $\mu\simeq 0.9\times 10^{-21}$ eV this threshold is $M_{\rm crit}\simeq 8.5\times 10^{9}M_{\odot}$. Hence, dynamical friction acts for $M_{\rm BH}\gtrsim M_{\rm crit}$ in the mass integral range $10^{5}-10^{10}M_{\odot}$. We denote with band parameter range extending to conservative $170\times\Gamma_{211}^{-1}<\tau_{\mathrm{U}}$. This includes contributions from partially populated boson clouds. Significant fraction of SMBHs is expect to have large spin Volonteri:2004cf . We consider a scenario where initial maximal spin is possessed by all SMBHs, although our results can be readily linearly rescaled for smaller fraction.
In Fig. 4 we show that the GA cloud effects of our scenario can suppress the PTA strain by $\mathcal{O}(10-30)\%$ relative to the vacuum prediction, improving agreement with the observed PTA amplitude. Performing a $\chi^{2}$ statistical fit, we find that predicted strain in GA cloud scenario gives $\chi^{2}=16.5$ (blue solid - all SMBHs with maximal spin), $\chi^{2}=17.5$ (blue dashed - conservative superradiance timescales), while predicted strain in GW scenario gives $\chi^{2}=18.8$ (black solid - general relativity prediction). For completeness we also note that saturated GA clouds could continuously radiate in GWs Brito:2017wnc . This lies in $\sim 0.1$ microhertz frequencies for $10^{-21}\,\text{eV}$ ULBs and contributes negligibly to the nanohertz background.
Since the superradiant cloud drag becomes ineffective once the SMBH orbit shrinks inside its Bohr radius $r_{0}$, GW background exhibits a turnover at
$f_{b}\simeq 3.2~{}\mathrm{nHz}~{}\Big{(}\frac{\mu}{10^{-21}\,\mathrm{eV}}\Big{)}^{3}\Big{(}\frac{M_{\rm BH}}{10^{10}\,M_{\odot}}\Big{)}^{2}~{}.$ (9)
This is a distinct testable feature of our scenario, distinguishing it from other proposals such as those based on SMBH DM spikes (e.g. 2024ApJ…962L..40C ; Alonso-Alvarez:2024gdz ). Detecting or ruling out the predicted turnover with upcoming precision PTA observations would thus isolate the dynamical imprint of superradiant clouds and pin down ULB mass scales, with no assumption on DM abundance.
If ULBs constitute a significant fraction of the DM, they can also form galactic-scale solitonic cores whose dynamical friction can also impact GW observations Aghaie:2023lan . In the Supplemental Material we demonstrate that including these soliton effects for ULB masses around $\mu\sim 10^{-21}$ eV can alter the PTA-observable strain and be distinguishable. Importantly, however, besides distinct signatures in contrast to our scenario ULB soliton-driven effects are in tension with existing constraints on ULB DM such as Lyman-$\alpha$ Antypas:2022asj .
Conclusions.— ULB scalars with masses $\mu\sim 10^{-22}-10^{-21}$ eV can efficiently form superradiant clouds around rotating SMBHs and we show they can significantly impact binary dynamics. The resulting ULB dynamical friction can drive SMBH binaries through the long-standing final parsec evolution bottleneck, where SMBH coalescence can stall, in $\lesssim 1$ Gyr timescales. This induces suppression in nanohertz stochastic GW background from SMBH mergers and imprints a distinctive spectral turnover. Considering century-long observations of OJ287 system we set novel constraints on ULB masses $\mu\simeq(8.5-22)\times 10^{-22}$ eV that are completely independent of DM assumptions and show that ULB drag could also reconcile any future confirmations of the debated excess in orbital-decay power. Forthcoming PTA data and high-precision orbital monitoring of BH systems will decisively test our scenario. Linking particle physics, BH astrophysics, and GW cosmology our framework opens an interdisciplinary avenue for probing new fundamental fields with observations.
Acknowledgments.— We thank Masha Baryakhtar and Yong Tang for comments. H.Y.Z. would like to thank Sida Lu and Yi Wang for the fruitful discussion on final-parsec problem. Q.D. and H.Y.Z. were supported by IBS under the project code, IBS-R018-D3. M.H. was supported by IBS under the project code, IBS-R018-D1. V.T. acknowledges support by the World Premier International Research Center Initiative (WPI), MEXT, Japan and JSPS KAKENHI grant No. 23K13109.
SUPPLEMENTAL MATERIAL
Superradiant Bosons Driving Supermassive Black Hole Mergers
Qianhang Ding, Minxi He, Volodymyr Takhistov, Hui-Yu Zhu
Here we provide additional computational details of superradiant GA cloud analysis, SMBH merger rates and GW signatures of solitons.
I Scalar-Field Superradiant Gravitational Atom
A scalar field of mass $\mu$ in the Kerr background of a rotating BH of mass $M_{\rm BH}$ and dimensionless spin $\chi$ obeys the Klein-Gordon equation
$\Big{(}\Box_{\rm Kerr}-\mu^{2}\Big{)}\Phi=0~{},$ (S1)
where $\Box_{\rm Kerr}$ is the d’Alembertian operator associated with the Kerr metric. With appropriate boundary conditions, Eq. (S1) separates into Teukolsky radial and angular equations describing perturbations of a Kerr BH. This can be solved numerically with Leaver’s method Teukolsky:1973ha ; Leaver:1985ax . We consider analytical solutions in the non-relativistic limit.
For the gravitational fine-structure constant $\alpha\equiv\mu M_{\rm BH}\ll 1$ one can factor out the fast oscillating phase
$\Phi(t,\mathbf{x})=\frac{e^{-i\mu t}}{\sqrt{2\mu}}\,\psi(t,\mathbf{x})+\mathrm{c.c.}~{},$ (S2)
such that Eq. (S1) reduces at leading order in $\alpha$ to the hydrogen atom analogue Schrödinger-like equation describing the “gravitational atom (GA)”
$i\partial_{t}\psi=\left(-\frac{\nabla^{2}}{2\mu}-\frac{\alpha}{r}+{\cal O}(\alpha^{2})\right)\psi~{}.$ (S3)
Then, the bound GA states are labeled by quantum numbers $(n,\ell,m)$ with wavefunctions
$\psi_{n\ell m}=R_{n\ell}(r)\,Y_{\ell m}(\theta,\phi)\,e^{-i(\omega_{n\ell m}-\mu)t}~{},$ (S4)
where $R_{n\ell}$ denotes radial functions and $Y_{\ell m}$ spherical harmonics.
The GA eigenfunctions are described by complex eigenfrequencies $\omega_{n\ell m}=E_{n\ell m}+i\Gamma_{n\ell m}$ with real and imaginary components given as (e.g. Detweiler:1980uk ; Baumann:2018vus ; Baumann:2019eav )
$\begin{cases}E_{n\ell m}=\mu\Bigl{[}1-\dfrac{\alpha^{2}}{2n^{2}}-\dfrac{\alpha^{4}}{8n^{4}}-\dfrac{(3n-\ell-1)\alpha^{4}}{n^{4}(\ell+1/2)}+\dfrac{2\chi m\alpha^{5}}{n^{3}\ell(\ell+1/2)(\ell+1)}+{\cal O}(\alpha^{6})\Bigr{]},\\ \Gamma_{n\ell m}=2\,\tilde{r}_{+}\,C_{n\ell}\,g_{\ell m}\bigl{(}m\Omega_{\rm BH}-\omega_{n\ell m}\bigr{)}\,\alpha^{4\ell+5},\end{cases}$ (S5)
where $\tilde{r}_{+}$ denotes outer horizon of the BH, $C_{nl}$ are numerical coefficients, and $g_{lm}=g_{lm}(\chi,\alpha,\mu)$ are functions defined in e.g. Refs. Detweiler:1980uk ; Baumann:2018vus ; Baumann:2019eav . Exponential superradiant growth occurs when condition $m\Omega_{\rm BH}>\omega_{n\ell m}$ is satisfied, which happens first for the fastest-growing mode $|211\rangle$. Higher GA $\ell$-modes grow significantly more slowly with $\Gamma\propto\alpha^{4\ell+5}$ since $\alpha\ll 1$. Hence, the BH retains finite spin over cosmic times.
The cloud density $\rho_{nlm}=M_{nlm}|\psi_{nlm}|^{2}$ can be written as
$\rho_{nlm}(r,\theta)=\beta\,A_{nlm}(x,\theta)\dfrac{M_{\rm BH}}{r_{0}^{3}}$ (S6)
with $x\equiv r/r_{0}$, $r_{0}=(\mu\alpha)^{-1}$ being the Bohr radius, and $\beta\equiv M_{nlm}/M_{\rm BH}$. For the dominant states
$\rho_{211}(x,\theta)$ $=\frac{\beta}{64\pi}\dfrac{M_{\rm BH}}{r_{0}^{3}}x^{2}e^{-x}\sin^{2}\theta,$ (S7)
$\rho_{31-1}(x,\theta)$ $=\frac{\beta}{6561\pi}\dfrac{M_{\rm BH}}{r_{0}^{3}}(x^{2}-6x)^{2}e^{-2x/3}\sin^{2}\theta.$ (S8)
In Fig. S1 we display the corresponding profiles and enclosed cloud mass for parameters relevant to OJ287 ($\alpha=\beta=0.1$).
Figure S1: [Top] Density profiles of the $|211\rangle$ and $|31-1\rangle$ GA states for $\alpha=\beta=0.1$ around primary BH in OJ287 system. [Bottom] Ratio of enclosed superradiance cloud mass relative to SMBH $M_{\rm BH}$. The vertical dashed line mark the semi-major axis ($0.056$ pc) of OJ287. The dotted line corresponds to $50\,r_{s}$.
II Impact of Binary Companion on Gravitational Atoms
In BH binaries, boson cloud states could be affected by a companion BH. A companion of mass $M_{\ast}$ at an orbital separation $R_{\ast}(t)$ can perturb the boson cloud gravitationally. In the reference frame of central BH, considering $\vec{r}=\{r,\theta,\phi\}$ and keeping only the Newtonian contributions of the companion’s potential, the gravitational perturbation due to binary companion is
$V_{\ast}(t,\vec{r})=-\alpha\,q\sum_{\ell_{\ast}\geq 2}\sum_{|m_{\ast}|\leq\ell_{\ast}}\mathcal{E}_{\ell_{\ast}m_{\ast}}(\iota_{*},\varphi_{*})\,Y_{\ell_{\ast}m_{\ast}}(\theta,\phi)\,F_{\ell_{\ast}}(r)$ (S9)
where
$F_{\ell_{\ast}}(r)=\frac{r^{\ell_{\ast}}}{R_{\ast}^{\ell_{\ast}+1}}\,\Theta(R_{\ast}-r)+\frac{R_{\ast}^{\ell_{\ast}}}{r^{\ell_{\ast}+1}}\,\Theta(r-R_{\ast})$ (S10)
and $\Theta$ is the Heaviside step-function, $q=M_{\ast}/M_{\text{BH}}$ is the mass ratio of the companion and the GA. Here we have also introduced the tidal moments
$\mathcal{E}_{l_{\ast}m_{\ast}}=\frac{4\pi}{2l_{\ast}+1}Y^{\ast}_{l_{\ast}m_{\ast}}(\theta_{\ast},\phi_{\ast})~{}.$ (S11)
The angle $\theta_{\ast}$ represents the angle between the BH spin direction and the companion location, while $\phi_{\ast}$ represents the projection of the angular phase of the companion orbit with respect to the primary BH onto the BH equatorial plane. We note that in the Fermi normal reference frame the dipole contribution ($\ell_{\ast}=1$) vanishes Baumann:2018vus ; Baumann:2019ztm , although Tomaselli:2023ysb claims that a residual dipole might exist. Here, we focus on contributions starting with the quadrupole ($\ell_{\ast}=2$).
The companion perturbation can induce mixing of different GA eigenstates. Considering boson cloud states $|n\ell m\rangle$ and $|n^{\prime}\ell^{\prime}m^{\prime}\rangle$ the matrix element factorizes,
$\langle n^{\prime}\ell^{\prime}m^{\prime}|V_{\ast}|n\ell m\rangle$ $=(-1)^{m^{\prime}+1}\alpha q\sum_{\ell_{\ast},m_{\ast}}\mathcal{E}_{\ell_{\ast}m_{\ast}}\,\mathcal{G}_{-m^{\prime}m_{\ast}m}^{\ell^{\prime}\ell_{\ast}\ell}\,I_{r},$ (S12)
where $I_{r}$ represents the radial integral
$I_{r}=\,\int_{0}^{R^{\ast}}r^{2}{\rm{d}}rR^{\ast}_{n^{\prime}l^{\prime}}(r)R_{nl}(r)\frac{r^{l_{\ast}}}{R_{\ast}^{l_{\ast}+1}}+\int_{R^{\ast}}^{\infty}r^{2}{\rm{d}}rR^{\ast}_{n^{\prime}l^{\prime}}(r)R_{nl}(r)\frac{R_{\ast}^{l_{\ast}}}{r^{l_{\ast}+1}}~{}.$
Here, $\mathcal{G}^{l^{\prime}l_{\ast}l}_{-m^{\prime}m_{\ast}m}$ is the Gaunt integral
$\mathcal{G}^{l^{\prime}l_{\ast}l}_{-m^{\prime}m_{\ast}m}=\int{\rm{d}}\Omega Y_{l^{\prime}-m^{\prime}}(\Omega)Y_{l_{\ast}m_{\ast}}(\Omega)Y_{lm}(\Omega)~{}.$ (S13)
Requiring that Gaunt integral does not vanish imposes that the following selection rules hold $-m^{\prime}+m_{\ast}+m=0$, $\lvert\ell-\ell^{\prime}\rvert\leq\ell_{\ast}\leq\ell+\ell^{\prime}$, and $\ell+\ell_{\ast}+\ell^{\prime}=2p,~{}p\in\mathbb{Z}$.
With the Newtonian potential, we can construct the new Hamiltonian $H=H_{0}+V_{\ast}$ consisting of a free part
$H_{0}=-\frac{1}{2\mu}\partial^{2}_{\mathbf{r}}-\frac{\alpha}{r}+\mathcal{O}(\alpha^{2})~{},$ (S14)
and a tidal perturbation part $V_{\ast}$. The existence of $V_{\ast}$ leads to three phenomena: direct off-resonance coupling of eigenmodes, Landau-Zener (LZ) resonance transitions, and ionization.
For illustration, we consider two cloud states $|1\rangle$ and $|2\rangle$ mixed by the tidal potential $V_{\ast}$ of the companion. We define the state coupling as $\eta(t)=V_{21}\langle 2|V_{\ast}|1\rangle$ and take $m_{1,2}$ to be the azimuthal quantum numbers of the two levels. The Hamiltonian in the interaction basis is
$H=\begin{pmatrix}\omega_{1}+V_{11}&V_{12}\\[2.0pt] V_{21}&\omega_{2}+V_{22}\end{pmatrix}=\begin{pmatrix}\bar{E}_{1}(t)+i\Gamma_{1}&\eta^{*}(t)\\[2.0pt] \eta(t)&\bar{E}_{2}(t)+i\Gamma_{2}\end{pmatrix},$ (S15)
where $\bar{E}_{1,2}$ and $\Gamma_{1,2}$ are the unperturbed energy and growth rates, respectively.
Figure S2: Survival probability of the GA superradiant eigenmodes under successive resonances. The red curve marks $170\times\Gamma_{211}^{-1}=\tau_{\rm U}$, above which the $|211\rangle$ cloud can form within a Hubble time. [Top Left] $|211\rangle\to|31-1\rangle$ transition at $q=1$. [Top Right] $|31-1\rangle\to|62-2\rangle$ at $q=1$, with the black dashed line showing the 99% survival contour. [Bottom Left] $|211\rangle\to|31-1\rangle$ at $q=0.1$. The dashed curve is the 90% contour. [Bottom Right] $|211\rangle\to|31-1\rangle$ with the mass ratio of OJ287 binary, with the 99% survival contour.
Transforming to the frame co-rotating with the companion gives the leading modification of the superradiant rate of the growing level Tong:2022bbl ; Zhu:2024bqs . Taking $|1\rangle$ to represent the superradiant state such as $|211\rangle$ and $|2\rangle$ to represent the absorptive state such as $|21-1\rangle$, the instantaneous correction to superradiant rate of state $|1\rangle$ is
$\Delta\Gamma_{1}\simeq-\,\frac{\Gamma_{1}-\Gamma_{2}}{\bigl{[}(E_{1}-E_{2})-(m_{1}-m_{2})\dot{\varphi}_{\ast}(t)\bigr{]}^{2}}\,|\eta(t)|,$ (S16)
which is always negative and largest near resonance. For a general binary orbit and eccentricity, one should average over a period Fan:2023jjj
$\langle\Delta\Gamma_{1}\rangle=\frac{1}{T}\int_{0}^{2\pi}\Delta\Gamma_{1}\,\frac{d\varphi_{\ast}}{\dot{\varphi}_{\ast}},$ (S17)
which shows that the effect scales with $|\eta|$.
In the case of OJ287, the orbital angular momentum is nearly orthogonal to the BH spin, i.e. inclination angle $\iota\simeq 90^{\circ}$. This strongly suppresses the overlap $|\eta|$ and correction to the superradiant rate. We estimate $\langle\Delta\Gamma_{1}\rangle/\Gamma_{1}\ll 1$. Thus, off-resonant mixing can be safely neglected in our analysis.
The LZ resonance transitions constitute other prominent effects. We observe that Eq. (S16) develops a pole when the driving frequency $\dot{\varphi}_{\ast}$ sweeps through $\Omega_{r}=(E_{1}-E_{2})/(m_{1}-m_{2})$. Close to the crossing we linearize the chirp, $\dot{\varphi}_{\ast}(t)\simeq\dot{\varphi}_{\ast}(t_{0})+\gamma t$ where $\gamma=\gamma_{\text{GW}}+\gamma_{\text{DF}}+\gamma_{\text{BR}}$ takes into account the energy loss via GW, dynamical friction and back-reaction of resonance effect on the companion. Here $\gamma_{\text{GW}}$ can be calculated Eq. (2.28) in Baumann:2019ztm and $\gamma_{\text{DF}}$ can be estimated by $\langle P_{\text{DF}}\rangle/\langle P_{\text{GW}}\rangle\times\gamma_{\text{GW}}$. Since the back-reaction effect is subdominant, we ignore the $\gamma_{\text{BR}}$ in our calculation. and treat the two-level system with the standard LZ formalism Baumann:2019ztm ; Tomaselli:2024bdd . With initial populations $c_{1}(-\infty)=1$, $c_{2}(-\infty)=0$ the occupation after the passage is
$|c_{1}|^{2}=e^{-2\pi Z},\qquad|c_{2}|^{2}=1-e^{-2\pi Z},\qquad Z\equiv\frac{|\eta|^{2}}{\gamma},$ (S18)
where $Z$ is the usual LZ parameter.
For a saturated $|211\rangle$ cloud three LZ resonances are in principle relevant: $|211\rangle\to|31{-}1\rangle\to|62{-}2\rangle\to|200\rangle$ Baumann:2019ztm . The last state is rapidly absorbed, which we estimate as $\Gamma_{200}\sim 10^{-11}\,\mathrm{s^{-1}}$ for benchmark $\alpha=0.124$ as relevant for OJ287 primary source. Thus, the boson cloud will be depleted in roughly $\sim 3\times 10^{3}$ years once it transitions to the $|200\rangle$ state. However, during these resonances, the binary system is situated within the cloud. Full state depletion would require $Z\gg 1$ in the preceding evolution steps. In realistic binaries the external energy loss is dominated by GW radiation and the cloud’s dynamical friction, making $\gamma$ large and $Z\lesssim 10^{-3}$. Consequently each resonance is highly non-adiabatic and transfers at most a few per cent of the state. This significantly suppresses the LZ parameter $Z$, resulting in an incomplete resonance transition.
In Fig. S2 we display the survival probability of the $|211\rangle$ and $|31{-}1\rangle$ levels after the chain of Bohr-type LZ crossings, as a function of $(M_{\rm BH},\mu)$ and for several characteristic BH binary mass ratios $q$. For equal masses ($q=1$) the first crossing almost completely transfers state population from $|211\rangle$ to $|31{-}1\rangle$. However, $|31{-}1\rangle$ itself survives the subsequent crossing at the $>99\%$ level, and the density profiles of the two states differ by less than an order of magnitude as shown on Fig. S1. When the companion is lighter, e.g. $q=0.1$, more than $90\%$ of the cloud remains in $|211\rangle$ over the entire parameter space, and in the residual region even a complete transfer to $|31{-}1\rangle$ would leave that state $100\%$ intact considering $|31-1\rangle\to|62-2\rangle$ resonance. Considering OJ287 mass ratio $q\simeq 8\times 10^{-3}$ we find that survival of $|211\rangle$ state exceeds $99\%$ throughout its history. Hence, resonance depletion can be neglected in the analysis of OJ287.
We note that fine- and hyperfine-structure transitions require binary separations well above those reached by the OJ287 binary throughout its evolution history, for example tracing back to redshift $z\simeq 10$, and therefore never occur. Back-reaction of the few-percent state population that is shuffled between levels is negligible compared with the dominant dynamical-friction torque. Allowing for non-vanishing eccentricity or a $\iota\simeq 90^{\circ}$ inclination results in even greater survival fraction Tomaselli:2023ysb ; Tomaselli:2024bdd . We therefore consider for simplicity with circular coplanar orbit in the main text.
III Dynamical Friction in Boson Cloud
Parameters Value Ref.
Primary SMBH mass $(10^{10}M_{\odot})$ $1.8348\pm 0.0008$ Dey:2018mjg
Secondary SMBH mass $(10^{8}M_{\odot})$ $1.5013\pm 0.0025$ Dey:2018mjg
Eccentricity $e$ $0.657\pm 0.001$ Dey:2018mjg
Precession angle per period $\Delta\vartheta$ $38.62^{\circ}\pm 0.01^{\circ}$ Dey:2018mjg
Orbital period $T$ (yr) $12.067\pm 0.007$ Dey:2018mjg
Orbital period decay rate $\dot{T}$ $-0.00099\pm 0.00006$ Dey:2018mjg
Spin of the primary SMBH $\chi_{P_{2}}$ $0.381\pm 0.004$ Dey:2018mjg
Spin of the primary SMBH $\chi_{P_{1}}$ $0.313\pm 0.01$ Valtonen:2016awd
Semi-major axis $a$ (pc) $0.056$ 1997ApJ…481L…5V
Inclination angle of orbit plane $90^{\circ}$ Valtonen:2013spa
Table S1: Parameters characterizing OJ287 BH system.
A compact object of mass $M_{\ast}$ traversing a boson cloud of local density $\rho$, experiences a gravitational drag force acting as dynamical friction 1943ApJ….97..255C , which in the non-relativistic, wave-like regime reads Hui:2016ltb
$F_{\rm DF}=\frac{4\pi\,M_{\ast}^{2}\,\rho}{v^{2}}\,C_{\Lambda}(\xi,k\,r_{\Lambda})\,,$ (S19)
where $v$ is the speed relative to the boson cloud that equals $\alpha\sqrt{(1+q)/x_{\ast}}$ in circular orbit and $\alpha\sqrt{(1+q)(2/x_{\ast}-r_{0}/a)}$ in elliptic orbit with $x_{\ast}=R_{\ast}/r_{0}$, $\xi=M\mu/v$, $k=\mu v$, and $r_{\Lambda}=\min(r_{\rm orb},r_{0})$ with $r_{0}$ being the cloud’s Bohr radius and $r_{\rm orb}$ the orbit size. In the $\xi\ll 1$ limit the Coulomb logarithm becomes
$C_{\Lambda}(kr)=\mathrm{Cin}(2kr)+\frac{\sin 2kr}{2kr}-1~{},$ (S20)
with
$\mathrm{Cin}(z)=\int_{0}^{z}(1-\cos t)\frac{\mathrm{d}t}{t}~{}.$ (S21)
Writing the instantaneous frictional power as $P_{\rm DF}=F_{\rm DF}\,v$, and $q=M_{\ast}/M_{\rm BH}$ being the binary mass ratio, one finds
$P_{\rm DF}=4\pi\,\frac{q^{2}M_{\rm BH}^{2}}{\alpha\sqrt{(1+q)(2/x_{\ast}-r_{0}/a)}}\,\rho(x_{\ast},\theta_{\ast})\,C_{\Lambda}\!\bigl{(}k\,r_{\Lambda}\bigr{)}.$ (S22)
For eccentric orbits with semi-major axis $a$, eccentricity $e$, and eccentric anomaly $z$ with $r=a[1-e\cos z]$, the effective impact parameter $kr_{\Lambda}$ is given by Zhang:2019eid
$kr_{\Lambda}=\begin{cases}\sqrt{x_{p}}\,\dfrac{(1-e\cos z)^{3/2}}{\sqrt{1+e\cos z}}\,,&x_{p}\leq x_{97},\\[6.0pt] \dfrac{x_{97}}{\sqrt{x_{p}}}\,\dfrac{(1-e\cos z)^{3/2}}{\sqrt{1+e\cos z}}\,,&x_{p}>x_{97},\end{cases}$ (S23)
where $x_{p}=r_{p}/r_{0}$ is the periapsis in units of $r_{0}$, and $x_{97}$ encloses 97% of the cloud mass.
To estimate the impact of boson cloud-induced dynamical friction on SMBH binary evolution, we compute dynamical friction power averaged over one period. The orbit of SMBH binary can be described as
$R_{\ast}(\varphi_{\ast})=\frac{a(1-e^{2})}{1+e\cos{(1-\zeta)\varphi_{\ast}}}~{},$ (S24)
where $\zeta$ is the precession phase angle per period. When a secondary SMBH companion is traversing the GA boson cloud, it experiences position-dependent cloud density $\rho(R_{\ast},\theta_{\ast})$. The angular position $\theta_{\ast}$ depends on the orbital phase $\varphi_{\ast}$ and inclination angle $\iota$ as $\cos{\theta_{\ast}}=\cos{\varphi_{\ast}}\sin{\iota}$.
Then, the average of Eq. (S22) over one orbital period $T$ is
$\langle P_{\rm DF}\rangle=\frac{1}{T}\int_{0}^{2\pi}P_{\rm DF}(R_{\ast}(\varphi_{\ast}),\theta_{\ast}(\varphi_{\ast},\iota))\,\frac{\mathrm{d}t}{\mathrm{d}\varphi_{\ast}}\,\mathrm{d}\varphi_{\ast}~{}.$ (S25)
Here,
$\frac{\mathrm{d}\varphi_{\ast}}{\mathrm{d}t}=\frac{\sqrt{(1+q)M_{\rm BH}a(1-e^{2})}}{R_{\ast}(\varphi_{\ast})^{2}}~{}$ (S26)
and the orbital period $T$ can be calculated as
$T=\int_{0}^{2\pi}\frac{\mathrm{d}t}{\mathrm{d}\varphi_{\ast}}\mathrm{d}\varphi_{\ast}~{}.$ (S27)
We employ the resulting Eq. (S25) in the main text to compute the orbital decay rate due to ULB clouds.
To calculate the ULB-induced dynamical friction impact on OJ287 orbital decay rate, we use the orbital parameters of OJ287 system $a=0.056\,\text{pc}$, $e=0.657$, and $\zeta=\Delta\vartheta/2\pi\simeq 0.107$ as shown in Tab. S1. Considering these parameters in Eq. (S25), we obtain $\langle P_{\rm DF}\rangle$ as a function of boson mass $\mu$ as shown in Fig. 2. Considering $\langle P_{\rm tot}\rangle-\langle P_{\rm GW}\rangle$ allows us to set novel constraints on ULB masses or identify ULB masses that can explain a potential excess as described in the main text.
Meanwhile, this boson-induced dynamical friction can speed up the orbital shrinkage in Eq. (6), which can help solve the final parsec problem. In evaluating evolution time within the final parsec, we set a constant eccentricity in orbital shrinkage, since the evolution of eccentricity in dynamical friction dominated region is highly uncertain Quinlan:1996vp ; 2022MNRAS.511.4753G . Then we integrate Eq. (6) over evolving separations to obtain its evolution time as follows,
$\tau=\int\frac{qM_{\rm BH}^{2}}{\langle P_{\mathrm{DF}}\rangle+\langle P_{\mathrm{GW}}\rangle}\frac{1}{2a^{2}}\mathrm{d}a~{}.$ (S28)
[…]
Figure S3: Allowed SMBH mass and ULB mass parameter space where binary evolution time from $0.5\,\mathrm{pc}$ to $0.1\,\mathrm{pc}$ is shorter than $\tau_{\rm pc}<1\,\mathrm{Gyr}$, alleviating the final-parsec problem.
IV SMBH binary merger rate
We can estimate the SMBH binary merger rate by mapping galaxy mergers to SMBH coalescences, considering each galaxy hosts a central BH. For primary SMBH mass by $M_{\rm BH}$ and mass ratio by $q$, and the primary galaxy stellar mass by $M_{\rm g}$ with merging galaxy mass ratio $q_{\rm g}$, one has
$\frac{\mathrm{d}^{3}n}{\mathrm{d}z\,\mathrm{d}q\,\mathrm{d}M_{\rm BH}}=\frac{\mathrm{d}^{3}n_{\rm g}}{\mathrm{d}z\,\mathrm{d}q_{\rm g}\,\mathrm{d}M_{\rm g}}\,\frac{\mathrm{d}M_{\rm g}}{\mathrm{d}M_{\rm BH}}\,\frac{\mathrm{d}q_{\rm g}}{\mathrm{d}q}\,.$ (S29)
The galaxy merger rate is parametrized as NANOGrav:2023hfp ; Chen:2018znx
$\frac{\mathrm{d}^{3}n_{\rm g}}{\mathrm{d}z\,\mathrm{d}q_{\rm g}\,\mathrm{d}M_{\rm g}}=\frac{\Psi(M_{\rm g},z^{\prime})}{M_{\rm g}}\,\frac{\mathcal{P}(M_{\rm g},q_{\rm g},z^{\prime})}{T_{\rm g-g}(M_{\rm g},q_{\rm g},z^{\prime})}\,\frac{\mathrm{d}t}{\mathrm{d}z^{\prime}}\,,$ (S30)
where $\Psi$ is the galaxy stellar mass function, $\mathcal{P}$ the galaxy pair fraction, and $T_{\rm g-g}$ the galaxy merger timescale. The cosmic time-redshift relation $\mathrm{d}t/\mathrm{d}z$ in a flat $\Lambda$CDM cosmology is given by
$\frac{\mathrm{d}t}{\mathrm{d}z}=\frac{1}{(1+z)\,H(z)}\,,\quad H(z)=H_{0}\sqrt{\Omega_{\Lambda}+\Omega_{m}(1+z)^{3}}\,,$ (S31)
with Planck 2018 parameters $H_{0}=67.4\,$km s-1 Mpc-1, $\Omega_{m}=0.315$, $\Omega_{\Lambda}=0.685$ Planck:2018vyg .
Parameter Value Parameter Value
$\psi_{0}$ $-2.63$ $\mathcal{P}_{0}$ $0.033$
$\psi_{z}$ $-0.6$ $\beta_{p_{0}}$ $1$
$m_{\psi_{0}}$ $11.5$ $T_{0}$ (Gyr) $0.5$
$m_{\psi_{z}}$ $0.11$ $\beta_{t_{0}}$ $-0.5$
$s_{\psi_{0}}$ $-1.21$ $\gamma_{t_{0}}$ $-1$
$s_{\psi_{z}}$ $-0.03$ $f_{\rm g}$ $0.615$
Table S2: Parameter values used for evaluating the the SMBH binary merger rate.
Adopting the Schechter form for galaxy stellar mass function, we parametrize the relevant quantities as
$\Psi(M_{\rm g},z)$ $=\Psi_{0}\left(\frac{M_{\rm g}}{M_{\Psi}}\right)^{s_{\Psi}}\exp\left(-\frac{M_{\rm g}}{M_{\Psi}}\right)~{},$ (S32)
$\mathcal{P}(M_{\rm g},q_{\rm g},z)$ $=\mathcal{P}_{0}(1+z)^{\beta_{p_{0}}}~{},$ (S33)
$T_{\mathrm{g-g}}(M_{\rm g},q_{\rm g},z)$ $=T_{0}(1+z)^{\beta_{t_{0}}}q_{\rm g}^{\gamma_{t_{0}}}~{},$ (S34)
with
$\log_{10}\left(\frac{\Psi_{0}}{\mathrm{Mpc}^{-3}}\right)$ $=\psi_{0}+\psi_{z}z~{},$ (S35)
$\log_{10}\left(\frac{M_{\Psi}}{M_{\odot}}\right)$ $=m_{\psi_{0}}+m_{\psi_{z}}z~{},$ (S36)
$s_{\Psi}$ $=1+s_{\psi_{0}}+s_{\psi_{z}}z~{}.$ (S37)
[…]
$\log_{10}\left(\frac{M_{\rm BH}}{M_{\odot}}\right)=8.7+1.1\log_{10}\left(\frac{M_{\mathrm{bulge}}}{10^{11}M_{\odot}}\right)~{},$ (S38)
where the bulge mass can be expressed by the stellar mass of galaxy as Lang:2014cta ; Bluck:2014ila
$M_{\mathrm{bulge}}=f_{\rm g}M_{\rm g}~{}.$ (S39)
From Eqs. (S38) and (S39), the relation between the mass ratios $q_{\rm g}$ and $q$ can be found as
$q=q_{\rm g}^{1.1}~{}.$ (S40)
In Tab. S2 we display the numerical values of relevant parameters NANOGrav:2023hfp ; Alonso-Alvarez:2024gdz .
V Boson Solitons and Gravitational Waves
Figure S4: The characteristic strain of GW background produced via SMBH binaries with different impacts from ULB soliton cores and superradiant GA boson clouds. We consider GA boson clouds in $|211\rangle$ state, with critical SMBH mass satisfying $\Gamma_{211}=1/\tau_{\mathrm{U}}$ for green solid line and $\Gamma_{211}=170/\tau_{\mathrm{U}}$ for green dashed line. The shaded region depicts impact from superradiant boson clouds on SGWB, considering ULB masses in the range of $(0.9-3)\times 10^{-21}\,\text{eV}$. Impact of solitonic cores with ULB masses of $\mu=1.2,1.5\times 10^{-21}\,\mathrm{eV}$ is shown. The mass ratio $\varepsilon$ between soliton core and SMBH is set as $10^{-3}$. The observation data points come from NANOGrav NANOGrav:2023gor , EPTA EPTA:2023xxk , and PPTA Reardon:2023gzh are shown.
ULB DM can form compact solitonic cores around SMBHs. The resulting signatures imprinted on stochastic GW background allows restricting ULB masses in the range of $1.3\times 10^{-21}\,\mathrm{eV}$ to $1.4\times 10^{-20}\,\mathrm{eV}$ Aghaie:2023lan . We calculate dynamical friction $\langle P_{\mathrm{DF}}\rangle$ considering density profile of solitons described as Aghaie:2023lan
$\rho_{\mathrm{sol}}\simeq\rho_{0}e^{-2\alpha\mu r}$ (S41)
with
$\rho_{0}\simeq 5\times 10^{4}\varepsilon\left(\frac{M}{10^{8}\,M_{\odot}}\right)^{4}\left(\frac{\mu}{10^{-21}\,\mathrm{eV}}\right)^{6}\,M_{\odot}\,\mathrm{pc}^{-3}~{},$ (S42)
where $\varepsilon$ is the mass ratio between solitonic core and SMBH.
In Fig. S4 we display the GW strain for stochastic background from SMBH solitonic cores for various ULB DM masses considering population of SMBHs following Eq. (8) and SMBH mass distribution in the range of $10^{5}-10^{10}M_{\odot}$. We compare these results with dynamical friction effects from ULB GA boson clouds in $|211\rangle$ state considering SMBHs with mass greater than critical mass $M_{\mathrm{crit}}$ that satisfies condition $\Gamma_{211}=1/\tau_{\mathrm{U}}$. The dynamical friction effects from ULB-DM solitonic cores and from superradiant clouds imprint qualitatively different low-frequency turnovers in the PTA spectrum, allowing these two scenarios to be distinguished.